The first side of a triangle is 8 meters longer than twice the second side, The third side is 3 times the second side. The perimeter is 128 meters. I'm supposed to find the lengths of three sides... I've done and redone this problem, it doesn't seem to make sense...
help please O:-)
thank you
-a VERY greatful student in need

Answer by askmemath(368)

Let the 2nd side of the triangle = X
Then first side = 2X+8
And third side = 3X
Perimeter of a triangle is the sum of all its sides
i.e. X+2X + 8 + 3x = 6X+8
But we know this equals 128
i.e. 6x+8 = 128
Subtracting 8 on both sides
6X = 120
Dividing by 6 throughout
X = 20 which is the 2nd side
First side = 2X+ 8 = 2(20)+8 = 48
3rd side = 3X = 3(20) = 60